# Undergraduate Course Allocation

### v1

@@ Abstract
Keywords: market design, many-to-many matching, course allocation, approximate competitive equilibrium, random serial dictatorship, reserved seats, priorities
@@
Course allocation is a well-studied problem in the context of graduate business schools. Many institutions employ course bidding systems that require strategic decision-making from students, often leading to inefficient allocations (Sönmez and Ünver, 2010). Recently, several business schools replaced course bidding systems with the approximate competitive equilibrium from equal incomes mechanism (Budish, 2011; Budish et al., 2023), which seeks to alleviate these issues. This mechanism elicits student preferences over course schedules, assigns students almost-equal budgets of fake money, and internally determines an approximate competitive equilibrium, in which each student receives her most-preferred affordable course schedule subject to course prices. The competitive equilibrium is approximate due to a small market-clearing error, measured by the assignment of a course’s seats over and under its capacity. This mechanism treats all business school students equally, as students are randomly assigned budgets and face the same course prices.
@@
At undergraduate institutions, students are not treated equally, making course allocation a two-sided matching problem. Each course divides students into priority levels based on factors such as seniority and student department. In commonly-used course allocation mechanisms, students select courses in order of seniority (with ties broken randomly), and priorities based on student department are enforced by reserving seats in each course. Top US universities that closely follow this process include Princeton, Johns Hopkins, Duke, Vanderbilt, Washington University in St. Louis, Columbia, Notre Dame, and Carnegie Mellon. [^2] In general, there is some heterogeneity in how universities elect to assign seats to students. Some universities use priorities (e.g., Dartmouth) or reserves (e.g., Vanderbilt) to assign seats, whereas others delegate the decision to departments (e.g., Cornell). Some use two phases (e.g., UCLA), only allowing students to enroll in a limited number of courses in the first phase. Seniority can be based on academic year (e.g., Carnegie Mellon), number of earned credits (e.g., University of North Carolina), or time to graduation (e.g., Washington University in St. Louis). At some universities, few classes are over-enrolled (e.g., Caltech), whereas at others, the process is rather congested (e.g., UC Berkeley). In most cases, there are a small number of priority levels, rather than one (Budish, 2011) or complete strict priorities Roth (1984).
@@
In this paper, we approach the task of allocating course schedules to students in the context of a many-to-many two-sided matching problem with heterogeneous student preferences and weak course priorities. We propose a novel deterministic allocation mechanism, the Pseudo-Market with Priorities (PMP) mechanism, that uses fake money and competitive equilibrium to allocate course seats to students without transfers. This mechanism elicits student preferences for course schedules, assigns almost equal budgets to students, and finds an approximate competitive equilibrium allocation under priority-specific prices. The equilibrium prices respect course priorities by setting a cutoff priority level for each course, where students at the cutoff priority level pay a non-negative price, students above the cutoff pay nothing, and students below the cutoff cannot afford the course.
@@
The approximate competitive equilibrium from equal incomes mechanism is the special case of our mechanism in which no course prioritizes any student over any others. Our first result shows that, with arbitrary course priorities, the PMP mechanism maintains the same, small worst-case bound on the market-clearing error as Budish (2011). This worst-case bound is increasing in the dimension of the space of prices, which equals the number of courses in settings without priorities. Our result holds despite the expansion from one price per course to priority-specific prices in each course. The key insight of the proof is the introduction of a parameterization within the priority-specific price space. The parameterization simplifies the space of prices to a single value per course, enabling a simultaneous search for both the cutoff priority level and the corresponding price. This strategy simplifies the complexity of the problem and allows for the application of some previously established results.
@@
The PMP mechanism has several advantageous theoretical properties. Adjusting course capacities to accommodate over-enrolled courses, the PMP outcome ensures that no student can be assigned a preferred schedule without violating course priorities or capacity constraints (approximate stability). It is also not possible to reassign course seats among a group of students, benefitting some members and hurting none, while ensuring that the outcome respects course priorities to the same degree (approximate priority-constrained efficiency). Among similarly prioritized students, the PMP outcome limits envy; if a student prefers the schedule assigned to a student with the same or a lower priority level in each course, by removing one course from the preferred schedule, the student now prefers her own (schedule envy bounded by a single course). Finally, in large populations, the PMP mechanism is resistant to strategic manipulations by students (strategy-proof in the large).
@@
To better guide practitioners in the course allocation process, we compare the performance of the PMP mechanism with several alternative mechanisms. As a benchmark, we consider a mechanism, the Random Serial Dictatorship with course reserves (RSD) mechanism, which closely resembles what is used in practice. Students select courses in order of seniority (with ties broken randomly) and priorities based on student department and year of study are enforced by reserving seats in each course. [^3] In practice, courses set a large number of reserved seats at the beginning of the registration process. After students initially register, university departments relax the course reserve constraints to ensure full enrollment. We do not model this period or the process of relaxing reserved seats. Instead, we calculate the optimal course reserves that a university registrar should set at the beginning of the process if the number of reserved seats cannot be changed. We also consider the Deferred Acceptance mechanisms with single tie-breaking and multiple tie-breakings (denoted as DA-STB and DA-MTB, respectively). These mechanisms allocate courses to students by extending the Gale-Shapley algorithm (Gale and Shapley, 1962) to many-to-many matching, where priority ties are resolved by either a single tie-breaking rule or course-specific tie-breakings. The Deferred Acceptance mechanism has been successful in many practical applications, including matching applicants to residency programs, students to schools, and workers to job positions (see Roth, 2018). We analyze six additional mechanisms in Appendix D.
@@
We use university data on the final allocation of 6 02360236\,0236 023 students to $756$ courses from the Spring 2018 term at a private institution in the mid-Atlantic region. In total, 23 3692336923\,36923 369 course seats are occupied, with $11.2\%$ of courses enrolled at or above their reported capacities. We use student and course characteristics to estimate a model of student utilities and evaluate the performance of the PMP, RSD, DA-STB, and DA-MTB mechanisms in terms of each student’s preferred mechanism, the standard deviation of student utilities, and the percentage of students who experience schedule envy. The first metric compares each student’s satisfaction from her course schedule across the four mechanisms. The two other metrics compare the fairness of each mechanism’s outcome across students.
@@
Our results demonstrate that the PMP mechanism delivers a substantially fairer outcome than the RSD benchmark. Considerably fewer students - $8\%$ of the total population, or almost $500$ individuals - have envy toward a student of weakly lower priority in each course. No students have envy greater than a single course, as opposed to $3\%$ in the benchmark mechanism. Furthermore, the standard deviation of student utilities declines between $1.1\%$ and $5.6\%$ across years of study. Each of these results are weaker for the DA-STB and DA-MTB mechanisms, with the DA-MTB mechanism delivering a slightly better performance.
These fairness improvements come without sacrificing student satisfaction. For second- through fourth-year students, a larger percentage of students prefer their assigned schedule in the PMP mechanism than students who prefer the RSD benchmark. The difference between these groups ranges from $5.0\%$ and $8.9\%$ across years of study, with $59.2\%-79.1\%$ of students being indifferent. Among students receiving different schedules, the average increase in utility ranges from $5.1\%$ and $9.4\%$. For first-year students, $0.9\%$ more prefer the benchmark. This result is attributed to the gains for older students, who are assigned an average of $135$ more seats. Still, the PMP mechanism provides an $0.8\%$ average increase in their utility.
@@
In contrast to these papers, we only consider deterministic mechanisms and study a many-to-many matching problem. Budish (2011) introduces the idea of competitive markets that may only approximately clear. He proposes a deterministic mechanism, the approximate competitive equilibrium from equal incomes mechanism, which is approximately Pareto-efficient, strategy-proof in the large, and bounds student envy by a single course. This mechanism was successfully implemented at Wharton Business School and Columbia Business School (see Budish et al., 2017; Budish and Kessler, 2022). [^5] There are few other studies of real-world mechanisms. Budish and Cantillon (2012) draw on theory and field data to study a non-strategy proof course allocation mechanism at Harvard Business School. Bichler and Merting (2021) compare the performance of two random mechanisms at the Technical University of Munich. Rusznák, Biró and Fleiner (2021) analyze course allocation at a large university in Hungary. A key contribution of the present study is extending Budish’s approach to two-sided matching settings with weak course priorities, a central feature of the undergraduate course allocation problem.
@@
Two recent papers also analyze deterministic assignments in many-to-many matching settings. Nguyen and Vohra (2022) establish the existence of a competitive equilibrium when all agent preferences satisfy a novel property of geometric substitutes, which is a strict generalization of the gross-substitutes property (see Kelso and Crawford, 1982). When there is an upper limit on the number of goods $k$ an agent can acquire, they prove that there always exists an approximate competitive equilibrium with no good over-enrolled by more than $k-1$ units. This “good-by-good” bound dominates the aggregate market-clearing bound of Budish (2011) when preferences are close substitutes, whereas the aggregate market-clearing bound dominates when student preferences are close complements. Though our main theoretical results follow the aggregate market-clearing bound approach, we ensure in our simulations that each approximate competitive equilibrium allocation found has no courses that are over-enrolled by more than $k-1$ seats.
@@
Similarly, Lin et al. (2022) analyze deterministic allocation mechanisms through approximate competitive equilibria with a good-by-good bound. They consider complex feasibility constraints and prove the existence of an approximate competitive equilibrium that satisfies stability and fairness properties similar to the ones analyzed in this paper. We establish some additional properties of our mechanism, proving that it is strategy-proof in the large and results in an outcome that is approximately priority-constrained efficient. With simulations on randomly generated data, they compare their approach to draft mechanisms in the context of reassigning sporting event season tickets to families. We address the problem of allocating courses to undergraduate students by estimating preferences from real-world university data and comparing our proposed mechanism to a version of the mechanism used at many top universities in the United States. In Appendix D, we also provide a comparison of our mechanism with the deferred-acceptance mechanism with single and multiple tie-breakings Gale and Shapley (1962), the deferred-acceptance mechanism with minority reserves Hafalir, Yenmez and Yildirim (2013), some alternative pseudo-market mechanisms Budish (2011), and a variant of the probabilistic serial mechanism Bogomolnaia and Moulin (2001). Many of these mechanisms have been studied in many-to-one matching settings, but their properties in many-to-many matching with indifferences are not well-understood.
@@
The remainder of the paper is organized as follows. We model undergraduate course allocation in Section 1 and investigate the properties of the PMP mechanism in Section 2. Our simulation results are presented in Section 3. We offer concluding remarks in Section 4. Appendices A and B present proofs, Appendix C details the student utility model calibration, and Appendices D and E provide additional simulation results and materials.
@@ 1 Environment
  • $Q=(q_{1},...,q_{M})$ is a vector of course capacities; each course $c\in{\cal C}$ has $q_{c}$ students.
  • $V=(\succsim_{1},....,\succsim_{N})$ is a vector of student preferences over course schedules: subsets of ${\cal C}$. Students typically consider only a subset of all possible course schedules due to factors such as course meeting times and prerequisites. Additionally, each student can take at most $k$ total course seats. We assume that these restrictions are incorporated into student preferences and that student preferences are strict. We also assume, for simplicity of exposition, that $21\leq k\leq M/$ 2. These restrictions still permit students to view courses as substitutes or complements.
  • =$\mathcal{R}=\{r_{s,c}\}_{s\in{\cal S},c\in{\cal C}}$ is a course priority structure. $r_{s,c}\in\{1,...,R\}$ specifies the level of priority of student $s$ in course $c$, with $R\geq 1$ and a larger number meaning a higher level of priority. Multiple students can share the same level of priority in a course.
@@
Two examples of course priority structures are year-specific and department-specific priorities. Universities often categorize students into four years of study based on semesters or credits completed. In addition, universities place students into departments, each of which offers a set of courses. In a year-specific priority structure, $R=4$, with $r_{s,c}$ equal to the year of study of student $s$ in every course $c$. In a department-specific priority structure, $R=2$, with $r_{s,c}=2$ if student $s$ and course $c$ are in the same department and $r_{s,c}=1$ otherwise. Year-specific priorities give an advantage to students closer to graduation in every course, whereas department-specific priorities give an advantage to students that require a given course to graduate over those that do not. We discuss hybrid priority structures that combine year-specific priorities and department-specific priorities in Section 3.
@@
In the next section, we analyze approximate market equilibria, in which some courses might have seats assigned slightly above capacity or empty demanded seats unfilled. We call such allocations having a market-clearing error. To account for the market-clearing error, we use an approximate version of stability.
@@
An allocation $x$ is approximately stable if it is stable in the environment with capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$ for each course $c\in{\cal C}$.
The relationship between stability and approximate stability is similar to the one between the notions of Pareto efficiency and approximate Pareto efficiency in Budish (2011). An allocation $y$ Pareto dominates an allocation $x$ if there is at least one student who strictly prefers her course schedule in $y$ and all other students weakly prefer their course schedules in $y$. Allocation $x$ is Pareto efficient if no allocation $y$ Pareto dominates $x$, and approximately Pareto efficient if no $y$ that Pareto dominates $x$ assigns weakly fewer seats in each course. Approximate Pareto efficiency is a meaningful notion of efficiency in environments where a market-clearing error is present.
@@
An allocation $x$ is approximately priority-constrained efficient if, in the environment with capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$ for each course $c\in{\cal C}$, for each allocation $y$ that Pareto dominates $x$, there is a course $c^{\prime}\in{\cal C}$ for which $y_{c^{\prime}}\nsucceq_{c^{\prime}}x_{c^{\prime}}$.
As with approximate stability, we account for a possible market-clearing error in allocation $x$ by adjusting the course capacities to $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$. Except for this adjustment, our definition coincides with the priority-constrained efficiency introduced by Schlegel and Mamageishvili (2020).
@@
This definition provides a criterion of fairness for students whose priorities are weakly ordered in the same way across all courses and minimizes envy between such students.
@@
Mechanisms that utilize market equilibria fail to be strategy-proof when students can influence course prices. A mechanism is strategy-proof in the large if truthful reporting is approximately optimal in an environment with a large number of students. For market equilibria, this amounts to being strategy-proof in a limit market in which each student regards the market prices as exogenous to her reported preferences (see Azevedo and Budish, 2019; Budish, 2011). [^7] He et al. (2018) considers the related notion of asymptotic incentive compatibility for random allocations. To avoid unnecessary heavy notation early in the paper, we introduce this concept formally in the proof of Theorem 5.

### v2

@@ Abstract
Keywords: market design, many-to-many matching, course allocation, approximate competitive equilibrium, random serial dictatorship, reserves, priorities
@@
Course allocation is a well-studied problem in the context of graduate business schools. Many institutions employ course bidding systems that require strategic decision-making from students, often leading to inefficient allocations (Sönmez and Ünver, 2010). Recently, several business schools replaced course bidding systems with the approximate competitive equilibrium from equal incomes (A-CEEI) mechanism (Budish, 2011; Budish et al., 2023), which seeks to alleviate these issues. This mechanism elicits student preferences over course schedules, assigns students almost-equal budgets of fake money, and internally determines an approximate competitive equilibrium, in which each student receives her most-preferred affordable course schedule subject to course prices. The competitive equilibrium is approximate due to a small market-clearing error, measured by the assignment of a course’s seats over and under its capacity. This mechanism treats all business school students equally, as students are randomly assigned budgets and face the same course prices.
@@
At undergraduate institutions, students are not treated equally, making course allocation a two-sided matching problem. Each course divides students into levels of priority based on student characteristics like year of study and department. In commonly-used course allocation mechanisms, students select courses in order of seniority (with ties broken randomly) and priorities based on departments are enforced by reserving seats in each course. Top US universities that closely follow this process include Princeton, Johns Hopkins, Duke, Vanderbilt, Washington University in St. Louis, Columbia, Notre Dame, and Carnegie Mellon. [^2] In general, there is some heterogeneity in how universities elect to assign seats to students. Some universities use priorities (e.g., Dartmouth) or reserves (e.g., Vanderbilt) to assign seats, whereas others delegate the decision to departments (e.g., Cornell). Some use two phases (e.g., UCLA), only allowing students to enroll in a limited number of courses in the first phase. Seniority can be based on academic year (e.g., Carnegie Mellon), number of earned credits (e.g., University of North Carolina), or time to graduation (e.g., Washington University in St. Louis). At some universities, few classes are over-enrolled (e.g., Caltech), whereas at others, the process is rather congested (e.g., UC Berkeley). In most cases, courses have a small number of priority levels, rather than one (Budish, 2011) or complete strict priorities Roth (1984).
@@
In this paper, we approach the task of allocating course schedules to students in the context of a many-to-many two-sided matching problem with heterogeneous student preferences and weak course priorities. We propose a novel deterministic allocation mechanism, the Pseudo-Market with Priorities (PMP) mechanism, that uses fake money and competitive equilibrium to allocate course seats to students without transfers. This mechanism elicits student preferences over course schedules, assigns almost-equal budgets to students, and finds an approximate competitive equilibrium allocation under priority-specific prices. These prices respect course priorities by setting a cutoff priority level for each course, where higher priority students obtain seats in the course for free, students at the cutoff level of priority face some price, and lower priority students cannot afford a seat.
@@
The A-CEEI mechanism is the special case of our mechanism in which no course prioritizes any student over any others. Our first result shows that, with arbitrary course priorities, the PMP mechanism maintains the same, small worst-case bound on the market-clearing error found in Budish (2011) for the A-CEEI mechanism. This bound is increasing in the dimension of the space of prices, which equals the number of courses in settings without priorities. Our result holds despite the expansion from one price per course to priority-specific prices in each course. The key insight of the proof is the introduction of a parameterization within the priority-specific price space. The parameterization simplifies the space of prices to a single value per course, enabling a simultaneous search for both the cutoff priority level and the corresponding price. This strategy simplifies the complexity of the problem and allows for the application of some previously established results.
@@
The PMP mechanism has several advantageous theoretical properties. Adjusting course capacities to accommodate over-enrolled courses, the PMP mechanism’s outcomes ensure that no student can be assigned a preferred schedule without violating course priorities or capacity constraints (approximate stability). It is also not possible to reassign course seats among a group of students, benefitting some members and hurting none, while ensuring that the outcome respects course priorities to the same degree (approximate priority-constrained efficiency). Among similarly prioritized students, its outcome limits envy; if a student prefers the schedule assigned to a student with the same or a lower priority level in each course, by removing one course from the preferred schedule, the student now prefers her own (schedule envy bounded by a single course). Finally, in large populations, the PMP mechanism is resistant to strategic manipulations by students (strategy-proof in the large).
@@
To better guide practitioners in the course allocation process, we compare the performance of the PMP mechanism with several alternative mechanisms. As a benchmark, we consider a mechanism, the Random Serial Dictatorship with course reserves (RSD) mechanism, which closely resembles what is used in practice. Students select courses in order of seniority (with ties broken randomly) and priorities based on student department and year of study are enforced by reserving seats in each course. [^3] In practice, at some universities, courses set a large number of reserved seats at the beginning of the registration process. After students initially register, university departments relax the course reserve constraints and admit students from waiting lists to ensure full enrollment. We do not model this later period of the registration process. Instead, we estimate the optimal course reserves that a university registrar should set at the beginning of the process if the number of reserved seats cannot be changed. We also consider the Deferred Acceptance mechanism with a single and with multiple tie-breakings (denoted DA-STB and DA-MTB, respectively). These mechanisms allocate courses to students by extending the Gale-Shapley algorithm (Gale and Shapley, 1962) to many-to-many matching, where priority ties are resolved by either a single tie-breaking rule or course-specific tie-breakings. The Deferred Acceptance mechanism has been successful in many practical applications, including matching applicants to residency programs, students to schools, and workers to job positions (see Roth, 2018). We analyze six additional mechanisms in Appendix D.
@@
We use university data on the final allocation of 6 0236\,023 students to $756$ courses from the Spring 2018 term at a private institution in the mid-Atlantic region. In total, 23 36923\,369 course seats are occupied, with $11.2\%$ of courses enrolled at or above their reported capacities. We use student and course characteristics to estimate a model of student utilities and evaluate the performance of the PMP, RSD, DA-STB, and DA-MTB mechanisms with a variety of measures of student satisfaction and allocation fairness.
@@
Our results demonstrate that the PMP mechanism delivers a substantially fairer outcome than the benchmark RSD mechanism and the two DA mechanisms. Considerably fewer students - $8.1\%$ of the total population, or almost $500$ students - prefer the course schedule assigned to a student of weakly lower priority in each course. In the PMP mechanism, no student’s envy is greater than a single course, as opposed to $3.3\%$, $2.5\%$, and $0.7\%$ of students in the RSD, DA-STB, and DA-MTB mechanisms, respectively. While the PMP, DA-STB, and DA-MTB mechanisms prevent lower-priority students from taking a seat that a higher-priority student desires, $16.2\%$ of students would benefit from taking a seat that is assigned to a student with strictly lower priority in the RSD mechanism.
These fairness improvements come without sacrificing student satisfaction. Only $31\%$ of students change schedules between the PMP and RSD mechanisms. Among these students, $300$ more students strictly prefer their outcome in the PMP mechanism. In contrast, both DA mechanisms result in approximately the same number of students who strictly prefer the mechanism and who strictly prefer the RSD benchmark. The main difference between the two DA mechanisms is that only $11.3\%$ of students change their schedules for DA-STB, while $40.3\%$ do so for DA-MTB. The PMP mechanism also improves on the RSD benchmark in mean utility for students all years of study, unlike both DA mechanisms.
@@
In contrast to these papers, we only consider deterministic mechanisms and study a many-to-many matching problem. Budish (2011) introduces the idea of competitive markets that may only approximately clear. He proposes a deterministic mechanism, the approximate competitive equilibrium from equal incomes mechanism, which is approximately Pareto-efficient, strategy-proof in the large, and bounds student envy by a single course. This mechanism was successfully implemented at Wharton Business School and Columbia Business School (see Budish et al., 2017; Budish and Kessler, 2022). [^5] There are few other studies of real-world course allocation mechanisms. Budish and Cantillon (2012) use field data to study a non-strategy proof course allocation mechanism at Harvard Business School. Bichler and Merting (2021) compare the performance of two random mechanisms at the Technical University of Munich. Rusznák, Biró and Fleiner (2021) analyze course allocation at a large university in Hungary. A key contribution of the present study is extending Budish’s approach to two-sided matching settings with weak course priorities, a central feature of the undergraduate course allocation problem.
@@
Two recent papers also analyze deterministic assignments in many-to-many matching settings. Nguyen and Vohra (2022) establish the existence of a competitive equilibrium when all agent preferences satisfy a novel geometric substitutes property, which is a strict generalization of the gross substitutes property (see Kelso and Crawford, 1982). When there is an upper limit on the number of goods $k$ an agent can acquire, they prove that there always exists an approximate competitive equilibrium with no good over-enrolled by more than k1k-1 units. This “good-by-good” bound dominates the aggregate market-clearing bound of Budish (2011) when preferences are close substitutes, whereas the aggregate market-clearing bound dominates when student preferences are close complements. Though our main theoretical results follow the aggregate market-clearing bound approach, we ensure in our simulations that each approximate competitive equilibrium allocation found has no courses that are over-enrolled by more than k1k-1 seats.
@@
Similarly, Lin et al. (2022) analyze deterministic allocation mechanisms through approximate competitive equilibria with a good-by-good bound. They consider complex feasibility constraints and prove the existence of an approximate competitive equilibrium that satisfies stability and fairness properties similar to the ones analyzed in this paper. We establish some additional properties of our mechanism, proving that it is strategy-proof in the large and results in an outcome that is approximately priority-constrained efficient. With simulations on randomly generated data, they compare their approach to draft mechanisms in the context of reassigning sporting event season tickets to families. We address the problem of allocating courses to undergraduate students by estimating preferences from real-world university data and comparing our proposed mechanism to a version of the mechanism used at many top universities in the United States. In Appendix D, we also provide a comparison of our mechanism with the deferred acceptance mechanism with single and multiple tie-breakings Gale and Shapley (1962), the deferred acceptance mechanism with minority reserves Hafalir, Yenmez and Yildirim (2013), some alternative pseudo-market mechanisms Budish (2011), and a variant of the probabilistic serial mechanism Bogomolnaia and Moulin (2001). Many of these mechanisms have been studied in many-to-one matching settings, but their properties in many-to-many matching with indifferences are not well-understood.
@@
The remainder of the paper is organized as follows. We model undergraduate course allocation in Section 1 and investigate the properties of the PMP mechanism in Section 2. In Section 3, we present simulation results using university data. We offer concluding remarks in Section 4. Proofs are in Appendices A and B. Details on the student utility model calibration and additional simulation results are in Appendices C, D, and E.
@@ 1 Environment
  • $Q=(q_{1},...,q_{M})$ is a vector of course capacities; each course $c\in{\cal C}$ has $q_{c}$ seats.
  • $V=(\succsim_{1},....,\succsim_{N})$ is a vector of student preferences over course schedules: subsets of $\cal C$}. Students typically only consider a subset of all possible course schedules due to factors such as course meeting times and prerequisites. Additionally, each student can take at most $k$ total course seats. We assume that these restrictions are incorporated into student preferences and that student preferences are strict. We also assume, for simplicity of exposition, that $1\leq k\leq M/2$. These restrictions still permit students to view courses as substitutes or complements.
  • $\mathcal{R}=\{r_{s,c}\}_{s\in{\cal S},c\in{\cal C}}$ is a course priority structure. $r_{s,c}\in\{1,...,R\}$ specifies the level of priority of student $s$ in course $c$, with $R\geq 1$ and a larger number meaning a higher level of priority. Multiple students can share the same level of priority in a course.
@@
Universities often categorize students into four years of study based on semesters or credits completed. In addition, universities place students into departments, each of which offers a set of courses. These practices illustrate two examples of course priorities: year-specific and department-specific priorities. In a year-specific priority structure, $R=4$, with $r_{s,c}$ equal to the year of study of student $s$ in every course $c$. In a department-specific priority structure, $R=2$, with $r_{s,c}=2$ if student $s$ and course $c$ are in the same department and $r_{s,c}=1$ otherwise. Year-specific priorities give an advantage to students closer to graduation in every course, whereas department-specific priorities give an advantage to students that require a given course to graduate over those that do not. We discuss hybrid priority structures that combine year-specific priorities and department-specific priorities in Section 3.
@@
In the next section, we analyze approximate market equilibria, in which some courses might have seats assigned slightly above capacity or empty demanded seats unfilled. We say such allocations have a market-clearing error. To account for allocations with market-clearing errors, we use an approximate version of stability
@@
An allocation $x$ is approximately stable if it is stable with capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$ for each course $c\in{\cal C}$.
The relationship between stability and approximate stability is similar to the one between the notions of Pareto efficiency and approximate Pareto efficiency in Budish (2011). An allocation $y$ Pareto dominates an allocation $x$ if there is at least one student who strictly prefers her course schedule in $y$ and all other students weakly prefer their course schedules in $y$. Allocation $x$ is Pareto efficient if no allocation $y$ Pareto dominates $x$, and approximately Pareto efficient if no $y$ that Pareto dominates $x$ assigns weakly fewer seats in each course.
Approximate Pareto efficiency is a meaningful notion of efficiency in environments where a market-clearing error is present.
@@
An allocation $x$ is approximately priority-constrained efficient if, in an environment with capacity $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$ for each course $c\in{\cal C}$, for each allocation $y$ that Pareto dominates $x$, there is a course $c^{\prime}\in{\cal C}$ for which $y_{c^{\prime}}\nsucceq_{c^{\prime}}x_{c^{\prime}}$.
As with approximate stability, we account for a possible market-clearing error in the allocation $x$ by adjusting the course capacities to $q^{\prime}_{c}=\sum_{s\in{\cal S}}x_{s,c}$.
Except for this adjustment, our definition coincides with the priority-constrained efficiency introduced by Schlegel and Mamageishvili (2020).
@@
This criterion of fairness minimizes each student’s envy towards students with weakly lower priority in every course.
@@
Mechanisms that utilize market equilibria fail to be strategy-proof when students can influence course prices. A mechanism is strategy-proof in the large if truthful reporting is approximately optimal in an environment with a large number of students. For market equilibria, this amounts to being strategy-proof in a limit market in which each student regards the market prices as exogenous to her reported preferences (see Azevedo and Budish, 2019; Budish, 2011). [^7] He et al. (2018) and Liu and Pycia (2016) consider two related notions of asymptotic incentive compatibility and asymptotic strategy-proofness for cardinal and ordinal preferences, respectively. Nguyen, Teytelboym and Vardi (2025) recently introduced a stronger notion of uniform strategy-proofness. To avoid unnecessary heavy notation early in the paper, we introduce this concept formally in the proof of Theorem 5.